OpenStudy (anonymous):

Which value is a solution to the inequality j + 3 is less than or equal to 6? 9 7 4 2

OpenStudy (anonymous):

2 because 2+3=5 and it i less than 6

OpenStudy (anonymous):

\[j + 3 \le 6 \] Therefore \[j \le 6 - 3 \] \[j \le 3\] So 2 satisfies the inequality.

OpenStudy (anonymous):

@its perpetual why didnt you just use substitution method to find the answer? not that you answer isnt right it is just a simpler way?

OpenStudy (anonymous):

I think it's more clear laying it out that way.

OpenStudy (karatechopper):

2